OpenStudy (anonymous):

The linear combination method is applied to a system of equations as shown. 4(.25x + .5y = 3.75) → x + 2y = 15 (4x – 8y = 12) → x – 2y = 3 2x = 18 What is the solution of the system of equations? (1,2) (3,9) (5,5) (9,3)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (bohotness):

which one do yuo think is the answer?

OpenStudy (anonymous):

i believe a @bohotness

OpenStudy (bohotness):

okay how did you come up with you answers

jimthompson5910 (jim_thompson5910):

Let's test that claim if the solution was (x,y) = (1,2), then you'd plug it into each equation x + 2y = 15 1 + 2(2) = 15 ... plug in x = 1 and y = 2 1 + 4 = 15 5 = 15 ... this is false, so (1,2) is NOT the solution to this equation Since (1,2) is NOT the solution to the first equation, it is NOT the solution to the entire system of equations.

OpenStudy (anonymous):

nevemind i got the last one because of the reason that if you divide 18/2 which would equal the value of x @bohotness

OpenStudy (bohotness):

okay

OpenStudy (anonymous):

right @bohotness

jimthompson5910 (jim_thompson5910):

gabylovesu, test your potential answer back into the original equations like I did above